Let X be a continuous random variable with density function "f(x) = \begin{cases}
 \frac{x}{2} &\text{for } 0 \le x \le 2 \\
 0 &\text{, } otherwise
\end{cases}"

Find E[|X-E[X]|].

1
Expert's answer
2021-07-18T17:54:53-0400

"f(x) = \\begin{cases} \\frac{x}{2} &\\text{for } 0 \\le x \\le 2 \\\\ 0 &\\text{, } otherwise \\end{cases}\n\n\u200b"


"E[X]= \int_{-\infty}^\infty xf(x)dx"

Here,

"E[X]=\int _0^2x\times \frac{x}{2}dx=\int_0^2\frac{x^2}{2}dx=\dfrac{x^3}{6}]_0^2=\dfrac{4}{3}"


Now,

"E[|X-E[X]|]=E[|X-\frac{4}{3}|]"

So,


"E[|X-\frac{4}{3}|]=\int_0^2|x-\frac{4}{3}|f(x)dx=\int_0^2|x-\frac{4}{3}|(\frac{x}{2})dx=\int_0^2|\frac{x^2}{2}-\frac{2x}{3}|dx"


"\implies \dfrac{x^3}{6}]_0^2\ \ -\ \dfrac{x^2}{3}]_0^2=\dfrac{8}{6}-\dfrac{4}{3}=0"


Hence, "E[|X-E[X]|]=0"